Circular No. 7210

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999cr and 1999cs J. Maza, University of Chile; and M. Hamuy, University of Arizona,

report the discovery by R. Antezana of two supernovae on 20-min unfiltered T-Max films taken by L. Gonzalez with the University of Chile’s Maksutov telescope:

SN  1999 UT  α_{2000}  δ_{2000}  Mag.  Offset
1999cr Mar. 12.20 13^{h}20^{m}18.^{s}30 −20^{o}08′50 5 18.0  7′′ E, 4′′ N
1999cs Apr. 14.28 12 04 25.27 −20 06 02.0 17.5  7′′ S

R. C. Smith obtained a spectrogram of SN 1999cr in ESO 576-34 on Mar. 17, using the 4-m Blanco telescope at Cerro Tololo, showing this to be a fairly young type-II supernova at z ∼ 0.01. M. T. Ruiz and Hamuy obtained BVRI CCD images of SN 1999cs on Apr. 17.146 with the Cerro Tololo 0.9-m telescope; a spectrum obtained by M. Phillips on Apr. 23 with the Dupont 2.54-m telescope at Las Campanas Observatory shows SN 1999cs to be a type-Ia supernova near maximum light (z = 0.06).

SUPERNOVAE 1999ct, 1999cu, 1999cv A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise

Observatory Optical Transients Search (cf. IAUC 6917, 7055) their discovery of three apparent supernovae on unfiltered CCD images taken with the Wise Observatory 1-m telescope:

SN  1999 UT  α_{2000}  δ_{2000}  R  Offset
1999ct June 13 13^{h}13^{m}04.^{s}56 +46^{o}15′52 7 20.8 2′′.8 W, 1′′.0 N
1999cu June 13 13 13 15.42 +46 20 00.8 21.2  —
1999cv June 13 14 34 43.88 +54 44 26.5 20.8  —

SNe 1999ct and 1999cu are located in the field of the galaxy cluster Abell 1697 (z = 0.183), while SN 1999cv is located in the field of the galaxy cluster Abell 1936 (z = 0.139). SNe 1999cu and 1999cv are superimposed on compact host galaxies, so no offsets can be determined from the images. All three objects were discovered in frames obtained by Gal-Yam and A. Anter on June 13 and confirmed by Gal-Yam on June 15. SNe 1999ct and 1999cu were invisible in images obtained on Feb. 15, while SN 1999cv was invisible in images obtained in May 14 with the same telescope (limiting magnitude R ∼ 22.0).

1999 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green